OpenStudy (anonymous):

if f(x) = (2-x)^2 and the y-intercept of f(x) = 4/3, find the rule for f(x)

OpenStudy (zzr0ck3r):

rule?

OpenStudy (jdoe0001):

?

OpenStudy (zzr0ck3r):

Please rewrite the question exactly as asked. @a5657371

OpenStudy (anonymous):

that is the question

OpenStudy (zzr0ck3r):

the question does not make sense.

OpenStudy (anonymous):

okay, thanks for trying-(its a modelling and problem solving question)

OpenStudy (zzr0ck3r):

since \(f(x)=(2-x)^2\), and the y-intercept is the point \((x,y)\) when \(x = 0\) we see \(f(0)=(2-0)^2=4\) so the y-intercept is \((0,4)\). Having said that, the rest of the question makes no sense either.

OpenStudy (anonymous):

Thanks